### v1

# A net theoretic approach to homotopy theory

Abstract.
This paper uses a net-theoretic approach to convergence spaces, aimed to simplify the description of continuous convergence in order to apply it in problems concerning Homotopy Theory. We present methods for handling homotopies of limit spaces, define fundamental groupoids, and prove a generalized version of the Seifert-van Kampen Theorem for limit spaces.
Key words and phrases:
convergence spaces, nets, homotopy theory.
2020 Mathematics Subject Classification:
54A05,54A20,55Q05
Introduction
The term “convenient category of topological spaces”, introduced by Brown [3] and popularized by Steenrod [36], refers to any category $\mathcal{C}$ of topological spaces nice enough to address the needs of working algebraic topologists. In modern terminology [26], this consists in asking for $\mathcal{C}$ to be a cartesian closed, complete and cocomplete full replete subcategory of Top containing CW-complexes among its objects. In this sense, the primary example is the category of compactly generated spaces, also known as $k$-spaces. However, if one drops the requirement of being a subcategory of Top, in favor of something like the “nice categories” in [27], then we can discuss supercategories of Top in which the lack of “natural” function spaces is corrected, such as the categories of convergence spaces, the main setting of this paper.
Since Choquet [7] introduced the convergence structures known as pseudotopologies and pretopologies, there has been extensive study on the subject. This has resulted in several papers and books addressing the topic. Notable textbooks include those by Čech [13], Binz [2], Gähler [15], Schechter [35], Beattie & Butzmann [1], Preuss [32] and more recently, Nel [25], Dolecki & Mynard [11] and Dolecki [10]. Roughly speaking, a convergence space is a set equipped with a notion of convergence for filters (or nets) defined on it. In other words, it provides a way to relate these objects to points in the space, which are called their limit points. Topological spaces are prima facie convergence spaces, but the converse does not hold: as convergence structures are defined directly, there are some cases in which the structure cannot be induced by a topology. Unlike topological spaces, many classes of convergence spaces have natural function spaces, making the corresponding categories (theoretically) nice for algebraic topology. However, this approach seems to have received little attention.
Indeed, a quick search on ZBMath for papers classified simultaneously under MSC55-XX (Algebraic Topology) and MSC54A20 (Generalities in General Topology regarding convergence) or MSC54A05 (Topological spaces and generalizations) returns fewer than forty matches, indicating the limited attention this topic has received in the literature [^1] Among these matches, only five articles appear to be directly related to the subject: [8, 9] by Bogin and Demaria, [22] by Mielke,[21] by Lee and Min, and[31] by Preuss.. More recently, Dossena [12] and Rieser [33, 34] have worked explicitly with homotopy theory in the context of convergence spaces. However, in personal communications with researchers in Algebraic Topology, convergence spaces seem to be largely unknown [^2] For instance: https://math.stackexchange.com/q/4746423.. The authors believe this is due to the extensive use of filters in the subject, which might have contributed to its limited adoption and exploration in “mainstream” Algebraic Topology research.
Although filters provide a robust setting for developing convergence theory, the notation can be cumbersome and somewhat unintuitive for newcomers. The typical use of filters often involves basic set-theoretic operations on families of sets of sets (and so on), making it easy to lose track of computations and the geometric intuition. As an alternative, we propose the use of nets, which, while not a novel approach, offers a more intuitive framework. Indeed, several authors, such as Čech [13], Schechter [35], Kelley [19], Katětov [18], Poppe [30], and Pearson [29], have already discussed convergence spaces in terms of nets. More recently, O’Brien et al. [28] carefully examined the topic, even addressing certain ZFC-related issues. However, the focus of these works is primarily on Analysis, not Algebraic Topology. Our goal is to extend this net-theoretic approach to convergence theory and explore its applications to problems in Algebraic Topology, with the beginnings of Homotopy Theory as our starting point.
The article is organized as follows. Sections 1 and 2 describe the basic setting of convergence spaces and nets used throughout the work. In Section 3, we discuss how to handle homotopies of limit spaces to define fundamental groupoids and fundamental groups, utilizing a Pasting Lemma for limit spaces – which seems to be new in the literature. With these concepts, we prove the groupoid version of the Seifert-Van Kampen Theorem for limit spaces in Section 4. Finally, the fifth section provides further comments and questions, concluding the article.
1. Nets (and filters) and convergence spaces
In 1922, Moore and Smith [24] introduced nets to generalize sequences, allowing for the handling of various types of limits in Analysis. In modern terminology, a net in a set $X$ is a function $\varphi\colon\mathbb{D}\to X$, where $\mathbb{D}$ is a directed set: this means that $\mathbb{D}$ has a binary relation $\preceq$ which is both reflexive and transitive (i.e., it is a preorder) such that for every $a,b\in\mathbb{D}$ there is $c\in\mathbb{D}$ for which $a,b\preceq c$. By a tail set of $\varphi$ we mean a subset like $\varphi[a^{\uparrow}]$ for some $a\in\mathbb{D}$, where $a^{\uparrow}\coloneqq\{b\in\mathbb{D}:b\succeq a\}$, thus $\varphi[a^{\uparrow}]=\{\varphi(b):b\succeq a\}$.
For a topological space $\left\langle X,\tau\right\rangle$, we say that a net $\varphi$ $\tau$-converges to a point $x\in X$ if every $\tau$-open set $O\subseteq X$ containing $x$ also contains a tail set of $\varphi$, i.e., there is $a\in\mathbb{D}$ such that $\varphi(b)\in O$ for every $b\succeq a$. We abbreviate this by writing $\varphi\to_{\tau}x$ or $\varphi(d)\to_{\tau}x$. It is clear that sequences are specific cases of nets, and the definition of convergent sequences in Analysis is a particular instance of the general definition of net convergence.
Although simple, the following well-known result helps to motivate the discussion that follows.
Proposition 1.1 (Folklore).
For topological spaces $\left\langle X,\tau\right\rangle$ and $\left\langle Y,\tau^{\prime}\right\rangle$, a function $f\colon X\to Y$ is continuous if and only if $f\circ\varphi$ $\tau^{\prime}$-converges to $f(x)$ for every $x\in X$ and every net $\varphi$ in $X$ such that $\varphi\to_{\tau}x$.
Proof.
[…]
Proposition 1.1 shows that convergent nets in a topological space completely determine its topology. Therefore, a natural step towards generalization is to consider abstract convergences, i.e., functions of the form $\textsc{NETS}\left(X\right)\to\wp(X)$, where $\textsc{NETS}\left(X\right)$ represents the “set” of all nets in $X$ and $\wp(X)$ is the power set of $X$. There is, however, a technical problem: if $X\neq\emptyset$, then $\textsc{NETS}\left(X\right)$ is not a set, but rather a proper class (cf. Jech [16, pp.5,6] and Schechter [35, pp.169]). This arises because it is relatively easy to “direct” a set [^3] Every function $Y\to X$ can be regarded as a net, since $Y$ can be well ordered., so $\textsc{NETS}\left(X\right)$ is too big to be an ordinary set.
To handle this, notice that the tail sets of the nets are what truly matter in the topological definition. So, for $\varphi,\psi\in\textsc{NETS}\left(X\right)$, let us say that $\varphi$ and $\psi$ are tail equivalent if every tail set of $\varphi$ contains a tail set of $\psi$ and vice versa. This relation, despite potential ZFC complications, defines an equivalence relation on $\textsc{NETS}\left(X\right)$, where the equivalence classes correspond precisely to the filters on $X$.
Filters were introduced by Henri Cartan [6] in 1937 and popularized by Bourbaki. By a filter on a set $X$ we mean a family $\mathcal{F}$ of subsets of $X$ closed under finite intersections and closed upwards; it is proper if $\emptyset\notin\mathcal{F}$. We denote the family of proper filters in $X$ by $\textsc{Fil}^{*}\left(X\right)$. For a topological space $\left\langle X,\tau\right\rangle$, the family $\mathcal{N}_{x}$ of $\tau$-neighborhoods of $x$ is a typical example of filter, which is used to define topological convergence in terms of filters: a proper filter $\mathcal{F}$ on $X$ is said to $\tau$-converges to $x$ if $\mathcal{N}_{x}\subseteq\mathcal{F}$, what we abbreviate with $\mathcal{F}\to_{\tau}x$.
Both notions of convergence are related. First, notice that a net $\varphi$ in $X$ induces a filter $\varphi^{\uparrow}$ in a very natural way: we simply let $\varphi^{\uparrow}$ be the family of those subsets of $X$ containing a tail set of the net $\varphi$. This yields a correspondence
$(\cdot)^{\uparrow}\colon\textsc{NETS}\left(X\right)\to\textsc{Fil}^{*}\left(X\right)$
such that $\varphi\to_{\tau}x$ if and only if $\varphi^{\uparrow}\to_{\tau}x$. Notice that two nets are tail equivalent if and only if they induce the same filters.
On the other hand, if we could find a “right inverse”
$\Gamma\colon\textsc{Fil}^{*}(X)\to\textsc{NETS}\left(X\right)$
then we would have $\Gamma(\mathcal{F})\to_{\tau}x$ if and only if $\left(\Gamma(\mathcal{F})\right)^{\uparrow}=\mathcal{F}\to_{\tau}x$. There is, indeed, such a function, as shown by Bruns and Schmidt [5]. Here is how it works. Given a filter $\mathcal{F}\in\textsc{Fil}^{*}\left(X\right)$, define $\mathbb{D}_{\mathcal{F}}\coloneqq\left\{\left\langle x,F\right\rangle\in X\times\mathcal{F}:x\in F\right\}$, preordered by the relation $\leq$ such that $\left\langle x,F\right\rangle\leq\left\langle y,G\right\rangle$ if and only if $G\subseteq F$, which makes $\left\langle\mathbb{D}_{\mathcal{F}},\leq\right\rangle$ a directed set. In this setup, the net $\Gamma(\mathcal{F})\colon\mathbb{D}_{\mathcal{F}}\to X$, defined by projection onto the first coordinate, has the members of $\mathcal{F}$ as its tail sets, ensuring that $\left(\Gamma(\mathcal{F})\right)^{\uparrow}=\mathcal{F}$.
The correspondences $(\cdot)^{\uparrow}$ and $\Gamma$ allow us to freely switch between nets and filters when discussing convergence in topological contexts. This flexibility extends to convergence spaces as well. Typically, a convergence on a set $X$ is defined as a function $L\colon\textsc{Fil}^{*}\left(X\right)\to\wp(X)$ that satisfies certain additional conditions. However, using $(\cdot)^{\uparrow}$, we can extend $L$ to a “function” $L\colon\textsc{NETS}\left(X\right)\to\wp(X)$ compatible with tail equivalence. Thus, the task becomes one of translating these additional conditions from the language of filters to that of nets.
In this work, a preconvergence on a set $X$ is a “function” $L:\textsc{NETS}\left(X\right)\to\wp(X)$ such that $L(\varphi)=L(\psi)$ whenever $\varphi^{\uparrow}=\psi^{\uparrow}$. As usual, we write $\varphi\to_{L}x$ instead of $x\in L(\varphi)$, and say that $x$ is an $L$-limit of $\varphi$, or that $\varphi$ $L$-converges to $x$, etc. The pair $\left\langle X,L\right\rangle$ is referred to as a preconvergence space [^4] The definition presented here aligns with Schechter’s [35], albeit Schechter refers to them as “convergence spaces”. However, terminology in this area remains inconsistent: for instance, Dolecki & Mynard [11] use “preconvergences” for what we call “isotone preconvergences”; similarly, Beattie &  Butzmann [1] and Nel [25] refer to “convergence spaces”, which Dolecki names “finitely stable spaces”, while we follow Preuss’s [32] terminology “limit spaces”. At least our use of “convergence spaces” follows Dolecki & Mynard’s terminology..
Remark 1.1.
As $\varphi^{\uparrow}=\left(\varphi(d):d\succeq D\right)^{\uparrow}$ for every $D\in\mathbb{D}$, we might always restrict a net to some of its tail sets without losing any of its limit points.
A function ${f}{\colon}X\to Y$ between preconvergence spaces $\left\langle X,L\right\rangle$ and $\left\langle Y,L^{\prime}\right\rangle$ is continuous if $f\left[L(\varphi)\right]\subseteq L^{\prime}(f\circ\varphi)$ for every $\varphi\in\textsc{NETS}\left(X\right)$, i.e., $f\circ\varphi\to_{L^{\prime}}f(x)$ whenever $\varphi\to_{L}x$. We denote by $\operatorname{C}\left(X,Y\right)$ the set of continuous functions from $X$ to $Y$. This data naturally defines a category, denoted by PrConv.
Now, we say a preconvergence $L$ on $X$ is
  1. (1)
    centered if constant nets $L$-converge to their corresponding constants,
  2. (2)
    isotone if $L(\varphi)\subseteq L(\psi)$ whenever $\varphi^{\uparrow}\subseteq\psi^{\uparrow}$,
  3. (3)
    stable if $L(\varphi)\cap L(\psi)\subseteq L(\rho)$,
where $\varphi$ and $\psi$ are nets on $X$, and $\rho$ is a mixing of $\varphi$ and $\psi$. The details of these conditions are briefly explained below.
Centerness should be clear: since we want constant functions to be continuous, we require constant nets to converge. Isotonicity, on the other hand, abstracts the behavior of subsequences in typical topological contexts. Notice that if $\left\langle y_{n}\right\rangle_{n}$ is a subsequence of $\left\langle x_{n}\right\rangle_{n}$, then $\left\langle x_{n}\right\rangle_{n}^{\uparrow}\subseteq\left\langle y_{n}\right\rangle_{n}^{\uparrow}$. Thus, in an isotone preconvergence, if $\left\langle x_{n}\right\rangle_{n}\to x$ and $\left\langle y_{n}\right\rangle_{n}$ is a subsequence of $\left\langle x_{n}\right\rangle_{n}$, we must have $\left\langle y_{n}\right\rangle_{n}\to x$ as well. With this in mind, we follow Schechter [35] an say that $\psi$ is a subnet of $\varphi$ if $\varphi^{\uparrow}\subseteq\psi^{\uparrow}$. Isotonicity, therefore, ensures that convergence is preserved for subnets. As Dolecki and Mynard [11], we say $\left\langle X,L\right\rangle$ is a convergence space if $L$ is both centered and isotone, and their category is denoted by Conv.
Finally, stability refers to the ability to mix two sequences converging to the same point to form a new sequence that also converges to that point. For instance, if $\left\langle x_{n}\right\rangle_{n}$ and $\left\langle y_{n}\right\rangle_{n}$ are sequences in a topological space converging to a point $z$, then $\left\langle z_{n}\right\rangle_{n}\to z$, where $z_{n}=x_{n}$ for even $n$ and $z_{n}=y_{n}$ for odd $n$. Mixing of nets generalizes this situation. Given two nets $\varphi,\psi\in\textsc{NETS}\left(X\right)$ on the same domain $\mathbb{D}$, a net $\rho\colon\mathbb{D}\to X$ is called a mixing of $\varphi$ and $\psi$ if there exists $D\in\mathbb{D}$ such that $\rho(d)\in\{\varphi(d),\psi(d)\}$ for every $d\succeq D$. Stability ensures that if $\rho$ is a mixing of the nets $\varphi$ and $\psi$, both converging to $x$, then $\rho$ also converges to $x$. A convergence spaces $\left\langle X,L\right\rangle$ is called a limit space if $L$ is stable, and the category of such spaces is denoted by Lim.
[…]
Remark 1.2.
Stability is also called finite depth in [11], which in filter terminology means that $L(\mathcal{F})\cap L(\mathcal{G})\subseteq L(\mathcal{F}\cap\mathcal{G})$ for every proper filters $\mathcal{F}$ and $\mathcal{G}$ on $X$. It is equivalent to the net version presented here if the preconvergence $L$ is isotone. We refer the reader to O’Brien et al. [28] for a proof, as well as other notions of mixings that we will not use in this work.
2. Convergential translations of topological notions
Given preconvergences $L$ and $L^{\prime}$ on a set $X$, we say $L^{\prime}$ is finer/stronger than $L$ (or $L$ is coarser/weaker than $L^{\prime}$), if $L^{\prime}(\varphi)\subseteq L(\varphi)$ for every $\varphi\in\textsc{NETS}\left(X\right)$, i.e, if $\varphi\to_{L^{\prime}}x$ implies $\varphi\to_{L}x$ for every $\varphi\in\textsc{NETS}\left(X\right)$ and $x\in X$, in which case we write $L\leq L^{\prime}$. This agrees with the usual topological terminologies: for topologies $\tau$ and $\tau^{\prime}$ on a set $X$, $\tau^{\prime}$ is finer than $\tau$ (i.e., $\tau\subseteq\tau^{\prime}$) if and only if the convergence $\to_{\tau^{\prime}}$ is finer than the convergence $\to_{\tau}$. This framework allows for the consideration of suprema and infima of preconvergences on a set $X$, making it possible to perform typical topological constructions, such as products, subspaces, coproducts and quotients. We summarize these constructions below.
For a family $\mathfrak{F}$ of preconvergences on a set $X$, its infimum and supremum in the family of all preconvergences on $X$ are realized by the preconvergences $\bigwedge\mathfrak{F}$ and $\bigvee\mathfrak{F}$, where
$\bigwedge\mathfrak{F}(\varphi)\coloneqq\bigcup_{L\in\mathfrak{F}}L(\varphi)\quad\text{and}\quad\bigvee\mathfrak{F}(\varphi)\coloneqq\begin{cases}\displaystyle\bigcap_{L\in\mathfrak{F}}L(\varphi)&\text{ if }\mathfrak{F}\neq\emptyset,\\ X&\text{ otherwise,}\end{cases}$
for every net $\varphi\in\textsc{NETS}\left(X\right)$. More precisely, $\varphi\to_{\bigwedge\mathfrak{F}}x$ if and only if there is a preconvergence $L\in\mathfrak{F}$ such that $\varphi\to_{L}x$, while for $\mathfrak{F}\neq\emptyset$, $\varphi\to_{\bigvee\mathfrak{F}}x$ if and only if $\varphi\to_{L}x$ for every preconvergence $L\in\mathfrak{F}$.
Suprema and infima are then used to define initial and final structures in the standard way. In fact, since these constructions rely on the lattice structure of the family of preconvergences on a set rather than the filter description of the convergences, all the well known results regarding these constructions remain valid [^5] Afterall, if anything, it is the description of convergence theory that is being changed, and not the theory itself.. In particular, the next lemma is a small fragment of a deeper result relating the so-called modifiers with the suprema of preconvergences in topological categories.
Lemma 2.1 (cf. [32]).
Let $\mathfrak{F}\neq\emptyset$ be a family of preconvergences on a set $X$. If $\left\langle X,L\right\rangle\in\textsc{Lim}$ for every $L\in\mathfrak{F}$, then $\left\langle X,\bigvee\mathfrak{F}\right\rangle\in\textsc{Lim}$.
Proof.
If $\varphi=\left\langle x\right\rangle_{d}$ is a constant net, then $\varphi\to_{L}x$ for every $L\in\mathfrak{F}$, hence $\varphi\to_{\bigvee\mathfrak{F}}x$. Assuming $\psi$ is a subnet of $\varphi$ such that $\varphi\to_{\bigvee\mathfrak{F}}x$, for every $L\in\mathfrak{F}$ we have $\varphi\to_{L}x$, then $\psi\to_{L}x$, thus implying $\psi\to_{\bigvee\mathfrak{F}}x$. The proof of stability is left to the reader. ∎
Remark 2.2.
[…]
For a family $\{\left\langle X_{i},L_{i}\right\rangle:i\in\mathcal{I}\}$ of preconvergence spaces, the cartesian product $\prod_{i\in\mathcal{I}}X_{i}$ carries the product preconvergence, in which a net $\varphi$ converges to $\left\langle x_{i}\right\rangle_{i\in\mathcal{I}}$ if and only if $\pi_{i}\circ\varphi\to_{L_{i}}x_{i}$ for each $i$. Similarly, a subset $X$ of a preconvergence space $\left\langle Y,L\right\rangle$ inherits the subspace preconvergence $L|_{X}$, in which a net $\psi\in\textsc{NETS}\left(X\right)$ $L|_{X}$-converges to $x\in X$ if and only if $\psi\to_{L}x$ in $Y$. By the previous lemma, it follows that $\prod_{i\in\mathcal{I}}X_{i}$ and $\left\langle X,L|_{X}\right\rangle$ are limit spaces provided all $\left\langle X_{i},L_{i}\right\rangle$ and $\left\langle Y,L\right\rangle$ are limit spaces.
Coproducts and quotients are defined dually. For a family $\{\left\langle X_{i},L_{i}\right\rangle:i\in\mathcal{I}\}$ of preconvergence spaces, their disjoint union $\coprod_{i\in\mathcal{I}}X_{i}$ carries the coproduct preconvergence, in which a net $\varphi$ converges to a point $w\in\coprod_{i\in\mathcal{I}}X_{i}$ if and only if there exist $j\in\mathcal{I}$, $x\in X_{j}$ and a net $\psi\in\textsc{NETS}\left(X_{j}\right)$ such that $\psi\to_{L_{j}}x$, $\iota_{j}(x)=w$ and $(\iota_{j}\circ\psi)^{\uparrow}\subseteq\varphi^{\uparrow}$ (here $\iota_{j}\colon X_{j}\to\coprod_{i\in\mathcal{I}}X_{i}$ is the canonical inclusion). Since the components $X_{i}$ are pairwise disjoint in the coproduct, it is straightforward to see that $\coprod_{i\in\mathcal{I}}X_{i}$ is a limit space provided that each $\left\langle X_{i},L_{i}\right\rangle$ is a limit space.
On the other hand, for a limit space $\left\langle X,L\right\rangle$ and a surjective function $f\colon X\to Y$, the quotient limit convergence on $Y$ induced by $f$ is defined such that a net $\varphi$ on $Y$ converges to $y$ if and only if there exist finitely many nets $\psi_{0},\dotso,\psi_{n}$ on $X$ and points $x_{0},\dotso,x_{n}\in X$ such that $\psi_{i}\to_{L}x_{i}$, $f(x_{i})=y$ for every $i\leq n$ and $\bigcap_{i\leq n}(f\circ\psi_{i})^{\uparrow}\subseteq\varphi^{\uparrow}$. The requirement for finitely many nets witnessing the convergence is done to ensure stability (cf. [1, 11, 32]).
All these constructions retain their standard universal properties, enabling further constructions like pushouts and pullbacks through the usual categorical recipes. As such, we omit a detailed discussion here and instead refer the reader to [32] or the more recent [23] for further guidance on these topics. We now turn our attention to translating other topological concepts.
Recall that a subset of a topological space is open if and only if it contains a tail set of every net converging to its points. With this in mind, we say a subset $O$ of a preconvergence space $\left\langle X,L\right\rangle$ is $L$-open if for every $\varphi\in\textsc{NETS}\left(X\right)$ for which $L(\varphi)\cap O\neq\emptyset$ there is $D\in\operatorname{dom}\left(\varphi\right)$ such that $\varphi[D^{\uparrow}]\subseteq O$. It is then not hard to see that $\tau_{L}\coloneqq\{O\subseteq X:O$ is $L$-open}}\}} is a topology on $X$. In fact, even more is true: $\tau_{L}$ is the strongest topology on $X$ such that the identity function $\left\langle X,L\right\rangle\to\left\langle X,\to_{\tau_{L}}\right\rangle$ is continuous and, by writing $\tau(X)$ to denote $\left\langle X,\to_{\tau_{L}}\right\rangle$, the correspondence $X\mapsto\tau(X)$ induces a functor [^6] Notice that if $f\colon X\to Y$ is continuous and $O\subseteq\tau(Y)$ is open, then $f^{-1}[O]\subseteq\tau(X)$ is open, for if $\varphi\to x$ in $X$, then $f\circ\varphi\to f(x)$ in $Y$, implying that $O$ contains a tail set of $f\circ\varphi$, hence $f^{-1}[O]$ contains a tail set of $\varphi$ as well. $\tau\colon\textsc{PrConv}\to\textsc{Top}$, usually called topological modification [1] or topologizer [10, 11]. Closed sets are defined as the complements of open sets.
The same strategy of adaption is followed by Dolecki and Mynard [11] to generalize the interior and closure operators in terms of filters, leading to the notions of inherences and adherences as their respective convergential counterparts. Under net terminology, the $L$-inherence of $S\subseteq X$ is the set of points in $X$ believing $S$ is an open set, i.e.,
$\operatorname{inh}_{L}\left(S\right)\coloneqq\left\{x\in X:\forall\varphi\in\textsc{NETS}\left(X\right)\quad\varphi\to_{L}x\Rightarrow S\in\varphi^{\uparrow}\right\},$
while the $L$-adherence of $S$ is $\operatorname{adh}_{L}\left(S\right)\coloneqq X\setminus\operatorname{inh}_{L}\left(X\setminus S\right)$. In particular, if $L$ is isotone, it can be showed that $x\in\operatorname{adh}_{L}\left(S\right)$ iff there is a net $\varphi\in\textsc{NETS}\left(S\right)$ such that $\varphi\to_{L}x$.
All these notions behave like their topological counterparts, in the following sense: given subsets $A,B,C,O,S\subseteq X$, it is straightforward to check that
  1. (1)
    $O\subseteq X$ is $L$-open iff $O\subseteq\operatorname{inh}_{L}\left(O\right)$,
  2. (2)
    $C\subseteq X$ is $L$-closed iff $\operatorname{adh}_{L}\left(C\right)\subseteq C$,
  3. (3)
    $\operatorname{adh}_{L}\left(\emptyset\right)=\emptyset$ and $\operatorname{inh}_{L}\left(X\right)=X$,
  4. (4)
    both $\operatorname{adh}_{L}\left(\cdot\right)$ and $\operatorname{inh}_{L}\left(\cdot\right)$ are $\subseteq$-increasing,
  5. (5)
    $\operatorname{adh}_{L}\left(A\cup B\right)=\operatorname{adh}_{L}\left(A\right)\cup\operatorname{adh}_{L}\left(B\right)$ and $\operatorname{inh}_{L}\left(A\cap B\right)=\operatorname{inh}_{L}\left(A\right)\cap\operatorname{inh}_{L}\left(B\right)$,
  6. (6)
    if $L$ is centered, then $S\subseteq\operatorname{adh}_{L}\left(S\right)$ and $\operatorname{inh}_{L}\left(S\right)\subseteq S$.
For centered spaces, one shows that $\operatorname{adh}_{L}\left(S\right)$ is $L$-closed for every $S\subseteq X$ if and only if the operator $\operatorname{adh}_{L}\left(\cdot\right)$ is idempotent [11, Proposition V.4.4], which in turns yields the classical characterization of topological pretopologies as the ones with idempotent adherences.
Remark 2.3.
As O’Brien et al. [28] do not mention inherences in their work, it is important to stress out that our definition is indeed equivalent to the one presented in [11]. Dolecki and Mynard say a point $x\in X$ is $L$-inherent to $S$ if $S\in\mathcal{F}$ for every proper filter $\mathcal{F}\in\textsc{Fil}^{*}\left(X\right)$ such that $\mathcal{F}\to_{L}x$. So, one just needs to replace $\mathcal{F}$ with a net $\Gamma(\mathcal{F})$ satisfying $\Gamma(\mathcal{F})^{\uparrow}=\mathcal{F}$.
Let us finally deal with compactness. We say a preconvergence space is compact if every net has a convergent subnet, what is simply the natural generalization of topological compactness in terms of nets [35]. To describe compactness for convergence spaces in term of covers, we need to replace open covers by something slightly more general.
We say a family $\mathcal{C}$ of subsets of $X$ is a local convergence system [^7] The usual terminology is “covering system”, as in [1]. at $x\in X$ if for all $\varphi\in\textsc{Nets}(X)$ such that $\varphi\to_{L}x$ there is $C\in\mathcal{C}$ containing a tail set of $\varphi$, i.e., such that $C\in\varphi^{\uparrow}$. We say that $\mathcal{C}$ is a convergence system if it is a local convergence system for each $x\in X$. With this terminology, it can be showed that a convergence space $\left\langle X,L\right\rangle$ is compact if and only if every convergence system has a finite subcover (cf. [1]).
In centered preconvergences, every convergence system is a cover. As for topological spaces, every convergence system $\mathcal{C}$ has a subcover $\mathcal{C}^{\prime}$ such that $\{\operatorname{int}{C}:C\in\mathcal{C^{\prime}}\}$ is an open cover. It also should be clear that for a continuous function $f\colon X\to Y$ between convergence spaces, $\{f^{-1}[C]:C\in\mathcal{C}\}$ is a convergence system for $X$ whenever $\mathcal{C}$ is a convergence system for $Y$. In this way, the following is an imediate corollary of the classical Lebesgue’s number lemma.
Proposition 2.1.
Let $H\colon K\to Z$ be a continuous function from a compact metric space $K$ to a convergence space $Z$. If $\mathcal{C}$ is a convergence system for $Z$, then there exists a real number $\delta>0$ such that every subset $A\subseteq K$ of diameter less than $\delta$ is contained in some $C\in\mathcal{C}$.
3. Homotopy and fundamental groupoids in limit spaces
The definition of homotopies between continuous functions remains traditional: given continuous functions $f,g\colon X\to Y$, where $X$ and $Y$ are limit spaces, a homotopy between $f$ and $g$ is a continuous function $H\colon[0,1]\times X\to Y$, with $[0,1]$ carrying its standard topological convergence, such that $H(0,x)=f(x)$ and $H(1,x)=g(x)$ for every $x\in X$. However, in our present context, we can naturally treat $H$ as a continuous path from $f$ to $g$ in $\operatorname{C}\left(X,Y\right)$, without imposing any additional conditions on $X$ and $Y$. All we have to do is to consider the continuous convergence structure over $\operatorname{C}\left(X,Y\right)$.
Given a net $\left\langle f_{d}\right\rangle_{d}$ on $\operatorname{C}\left(X,Y\right)$, we say that $\left\langle f_{d}\right\rangle_{d}$ converges continuously to $f\in\operatorname{C}\left(X,Y\right)$, denoted $f_{d}\to_{c}f$, if $\left\langle f_{d}(x_{a})\right\rangle_{d,a}$ converges to $f(x)$ in $X$ for every $x\in X$ and every net $\left\langle x_{a}\right\rangle_{a}$ in $X$ such that $x_{a}\to x$ (cf. [30, 28]). In filter terms, this is expressed by stating that a proper filter $\mathcal{F}$ on $\operatorname{C}\left(X,Y\right)$ converges continuously to $f$ if and only if $\mathcal{F}(\mathcal{G})\to f(x)$ for every $x\in X$ and every proper filter $\mathcal{G}$ on $X$ such that $\mathcal{G}\to x$, where $\mathcal{F}(\mathcal{G})$ is the filter generated by all sets of the form $F(G)\coloneqq\{f(x):f\in F\text{ and }x\in G\}$ with $F\in\mathcal{G}$ and $G\in\mathcal{G}$ (cf. [1, 10, 11, 32]). We refer the reader to [28] for a detailed discussion on the equivalence between these descriptions.
The net description of continuous convergence allows us to write really readable proofs of the main facts about $\left\langle\operatorname{C}\left(X,Y\right),\to_{c}\right\rangle$, in which the core ideas are not obscured by too many layers of sets.
  1. (1)
    The continuous convergence is a convergence to begin with, meaning that it is both centered and isotone. For centering, if the net $\left\langle f_{d}\right\rangle_{d}$ is constant, say $f_{d}=f$ for all $d$, and $x_{a}\to x$ in $X$, then $\left\langle f_{d}(x_{a})\right\rangle_{d,a}$ is tail equivalent to $\left\langle f(x_{a})\right\rangle_{a}$, which converges to $f(x)$ as $f$ is continuous. Regarding isotonicity, notice that $\left\langle f_{d}(x_{a})\right\rangle_{d,a}$ is a subnet of $\left\langle g_{e}(x_{a})\right\rangle_{e,a}$ whenever $\left\langle f_{d}\right\rangle_{d}$ is a subnet of $\left\langle g_{e}\right\rangle_{e}$, so $f_{d}\to_{c}g$ whenever $g_{e}\to_{c}g$, given that $Y$ is isotone by hypothesis.
  2. (2)
    $\left\langle\operatorname{C}\left(X,Y\right),\to_{c}\right\rangle$ is a limit space. For if $\left\langle h_{d}\right\rangle_{d}$ is a mixing of nets $\left\langle f_{d}\right\rangle_{d}$ and $\left\langle g_{d}\right\rangle_{d}$, both continuously converging to $h$, then $\left\langle h_{d}(x_{a})\right\rangle_{d,a}$ is still a mixing of $\left\langle f_{d}(x_{a})\right\rangle_{d,a}$ and $\left\langle g_{d}(x_{a})\right\rangle_{d,a}$ for every net $\left\langle x_{a}\right\rangle_{a}$ in $X$. This ensures that $h_{d}(x_{a})\to h(x)$ whenever $x_{a}\to x$, as $Y$ is a limit space.
  3. (3)
    $\left\langle\operatorname{C}\left(X,Y\right),\to_{c}\right\rangle$ is the exponential object $Y^{X}$ in the category Lim. This property highlights the significant advantage of convergence spaces over topological spaces. Since most proofs in the literature are written in filter terms, let us present a net version of it.
Proof of $(3)$.
First we show that the evaluation map $\operatorname{ev}\colon\operatorname{C}\left(X,Y\right)\times X\to Y$ is continuous. It goes as follows: if $\left\langle f_{d},x_{d}\right\rangle_{d}$ converges to $\left\langle f,x\right\rangle$ in $\operatorname{C}\left(X,Y\right)\times X$, then $f_{d}\to_{c}f$ in $\operatorname{C}\left(X,Y\right)$ and $x_{d}\to x$ in $X$, implying $\left\langle f_{d}(x_{d^{\prime}})\right\rangle_{d,d^{\prime}}\to f(x)$ in $Y$, hence $\operatorname{ev}(f_{d},x_{d})=f_{d}(x_{d})\to f(x)$ since $\left\langle f_{d}(x_{d})\right\rangle_{d}$ is a subnet of $\left\langle f_{d}(x_{d^{\prime}})\right\rangle_{d,d^{\prime}}$. Therefore, $\operatorname{ev}$ is continuous.
The next step is to show that for every limit space $Z$ and every continuous function $H\colon Z\times X\to Y$ there is a unique continuous function $\widehat{H}\colon Z\to\operatorname{C}\left(X,Y\right)$ such that the diagram below is commutative.
[…]
Notice that the commutativity of the diagram imposes that $\widehat{H}(z)\colon X\to Y$ maps each $x\in X$ to $H(z,x)$ as $z$ runs through $Z$, which is a continuous function because it is a composition of continuous functions. So, it remains to show that $\widehat{H}$ is continuous, what follows precisely by the definition of continuous convergence: if $z_{d}\to z$ in $Z$ and $x_{a}\to x$ in $X$, then $\left\langle z_{d},x_{a}\right\rangle\to\left\langle z,x\right\rangle$ in $Z\times X$, implying
$\widehat{H}(z_{d})(x_{a})=H(z_{d},x_{a})\to H(z,x)=\widehat{H}(z)(x),$
i.e., $\widehat{H}(z_{d})\to_{c}\widehat{H}(z)$. ∎
Remark 3.1 (Splitting and admissible convergences).
Adapting the standard terminology for topologies on functions spaces (cf. Engelking [14]), we say a convergence $\lambda$ on $\operatorname{C}\left(X,Y\right)$ is splitting if $\widehat{H}\colon Z\to\operatorname{C}\left(X,Y\right)$ is continuous whenever $H\colon Z\times X\to Y$ is continuous, and admissible if $\lambda$ makes the evaluation $\operatorname{ev}\colon\operatorname{C}\left(X,Y\right)\times X\to Y$ continuous. It follows by the proof above that the continuous convergence is both splitting and admissible. Conversely, it is not hard to show that a splitting admissible convergence must be the continuous convergence. In particular, $\to_{c}$ is the weakest convergence on $\operatorname{C}\left(X,Y\right)$ making the evaluation continuous.
Therefore, a classic homotopy $H\colon[0,1]\times X\to Y$ between continuous functions $f$ and $g$ gives rise to a (continuous) path $\widehat{H}\colon[0,1]\to\operatorname{C}\left(X,Y\right)$ from $f$ to $g$, in which $\operatorname{C}\left(X,Y\right)$ is endowed with the continuous convergence [^8] Unless otherwise stated, this is the only convergence structure we consider on $\operatorname{C}\left(X,Y\right)$.. This allows us to simplify some calculations when defining the fundamental groupoid of a limit space $X$: we will define a category $\Pi(X)$ whose objects are the points of $X$, such that an arrow $x\to y$ corresponds to a class of rel-homotopic paths from $x$ to $y$, essentially as one would do for topological spaces (cf. Brown [4] or Kammeyer [17]).
To fix terminology, a path in $X$ is a continuous function $\gamma\colon I\to X$, where $I$ stands for the interval $[0,1]$ with its standard topological convergence. We say $\gamma$ is a path from $x\in X$ to $y\in X$ if $\gamma(0)=x$ and $\gamma(1)=y$, which are the end points of the path. Two paths $\gamma,\gamma^{\prime}\colon I\to X$ with the same end points, say $x$ and $y$, are homotopic relative to its end points, abbreviated as rel-homotopic, if there is a path $H\colon I\to\operatorname{C}\left(I,X\right)$ from $\gamma$ to $\gamma^{\prime}$ such that $H(t)$ is a path from $x$ to $y$ for every $t\in I$.
Let $\gamma,\gamma^{\prime}\in\operatorname{C}\left(I,X\right)$ be two paths from $x$ to $y$. By writing $\gamma\simeq_{x,y}\gamma^{\prime}$ to indicate that $\gamma$ and $\gamma^{\prime}$ are rel-homotopic, the first issue is to show that $\simeq_{x,y}$ is an equivalence relation on the set of paths from $x$ to $y$. Reflexivity and symmetry are straightforward, but transitivity is more delicate, as it depends on a pasting lemma.
Lemma 3.2 (Pasting lemma).
Let $X$ and $Y$ be limit spaces and $f\colon X\to Y$ be a function. If $\mathcal{C}$ is a locally finite cover for $X$ by closed sets such that the restriction $f|_{C}$ is continuous for every $C\in\mathcal{C}$, then $f$ is continuous.
Proof.
For a point $x\in X$ and a net $\varphi$ in $X$ such that $\varphi\to x$, we need to show that $f\circ\varphi\to f(x)$ in $Y$. Let $O\subseteq X$ be an open set such that $x\in O$ and $\mathcal{S}\coloneqq\{C\in\mathcal{C}:C\cap O\neq\emptyset\}$ is finite. Since each $S\in\mathcal{S}$ is closed and $\mathcal{S}$ is finite, there is no loss of generality in assuming $x\in S$ for every $S\in\mathcal{S}$. Now, let $\varphi_{S}\colon\operatorname{dom}\left(\varphi\right)\to X$ be the net given by the rule
$\varphi_{S}(d)\coloneqq\begin{cases}\varphi(d)&\text{if }\varphi(d)\in S\\ \quad x&\text{if }\varphi(d)\not\in S\end{cases}.$
Since $X$ is a limit space [^9] In fact, it would be enough to ask $X$ to be a Kent space [32]., it follows that $\varphi_{S}\to x$. Given that $\varphi_{S}$ is a net in $S$ and $x\in S$, we get $\varphi_{S}\to x$ in the subspace $S$, so $f|_{S}\circ\varphi_{S}\to f|_{S}(x)$ in $Y$ due to the continuity of $f|_{S}$. As $Y$ is also a limit space, to conclude that $f\circ\varphi\to f(x)$, it suffices to find $A\in\operatorname{dom}\left(\varphi\right)$ such that $f(\varphi(\alpha))\in\{f|_{S}(\varphi_{S}(\alpha)):S\in\mathcal{S}\}$ for every $\alpha\succeq A$ (cf. Remark 1.1). Since $\varphi\to x$ and $x\in O$, there exists $A\in\operatorname{dom}\left(\varphi\right)$ such that $\varphi(\alpha)\in O$ for every $\alpha\succeq A$. Finally, notice that as $\mathcal{C}$ covers $X$, there is $C\in\mathcal{C}$ such that $\varphi(\alpha)\in C$, hence $C\cap O\neq\emptyset$, thus implying $C\in\mathcal{S}$ and $\varphi_{C}(\alpha)=\varphi(\alpha)$, so $f(\varphi(\alpha))=f|_{C}(\varphi_{S}(\alpha))$, as desired. ∎
Remark 3.3.
As far as we know, the formulation of the previous theorem is unknown in the literature: the version presented by Preuss [32] assumes $\mathcal{C}$ to be finite and $X$ to be pretopological; Dolecki and Mynard’s version [11] assume $Y$ to be pseudotopological; there is yet another version assuming both $X$ and $Y$ to be pseudotopological, due to Dossena [12]. On the other hand, Preuss [32] already showed that the result is false if $Y$ is not a limit space.
Therefore, for paths $\gamma,\gamma^{\prime}\in\operatorname{C}\left(I,Z\right)$ such that $\gamma(1)=\gamma^{\prime}(0)$, the usual gluing function $\gamma*\gamma^{\prime}\colon I\to Z$ given by
$\gamma*\gamma^{\prime}(t)\coloneqq\begin{cases}\gamma(2t)&\text{ if }t\leq\frac{1}{2}\\ \gamma^{\prime}(2t-1)&\text{ if }\frac{1}{2}\leq t\leq 1\end{cases}$
is such that $\gamma*\gamma^{\prime}$ is continuous in $\left[0,\frac{1}{2}\right]$ and $\left[\frac{1}{2},1\right]$, hence it is a path in $Z$ from $\gamma(0)$ to $\gamma^{\prime}(1)$. Since $\operatorname{C}\left(I,X\right)$ is a limit space whenever $X$ is a limit space, it follows that $\simeq_{x,y}$ defines an equivalence relation on the subset $\mathbb{P}[x,y]$ of $\operatorname{C}\left(I,X\right)$ whose members are the paths from $x$ to $y$: for paths $H,H^{\prime}\colon I\to\operatorname{C}\left(I,X\right)$ witnessing $\gamma\simeq_{x,y}\gamma^{\prime}$ and $\gamma^{\prime}\simeq_{x,y}\gamma^{\prime\prime}$ respectively, $H*H^{\prime}$ witnesses that $\gamma\simeq_{x,y}\gamma^{\prime\prime}$.
Proceeding with the construction of the category $\Pi(X)$:
  • objects are the points of $X$;
  • an arrow from $x$ to $y$ is a $\simeq_{x,y}$-equivalence class $[\gamma]$ of a path $\gamma\in\mathbb{P}[x,y]$.
Given arrows $[\gamma]\colon x\to y$ and $[\rho]\colon y\to z$, the composition $[\rho]\circ[\gamma]$ is the class $[\gamma*\rho]$. To see it is well defined, we take $\gamma^{\prime}$ and $\rho^{\prime}$ such that $\gamma\simeq_{x,y}\gamma^{\prime}$ and $\rho\simeq_{y,z}\rho^{\prime}$ and show that $\gamma*\rho\simeq_{x,z}\gamma^{\prime}*\rho^{\prime}$. First, for a path $G\colon I\to\operatorname{C}\left(I,X\right)$ witnessing $\gamma\simeq_{x,y}\gamma^{\prime}$, notice that $K\colon I\to\operatorname{C}\left(I,X\right)$, given by $K(t)\coloneqq G(t)*\rho$, is a function such that $K(t)\in\mathbb{P}[x,z]$ for which $K(0)=\gamma*\rho$ and $K(1)=\gamma^{\prime}*\rho$. So, in order to see that $K$ witnesses $\gamma*\rho\simeq_{x,z}\gamma^{\prime}*\rho$, it remains to verify its continuity, what follows easily once we notice that the function $I\times I\to X$, mapping $\left\langle t,s\right\rangle$ to $K(t)(s)$, is continuous when restricted to $I\times\left[0,\frac{1}{2}\right]$ and $I\times\left[\frac{1}{2},1\right]$. Similarly, one shows that $\gamma^{\prime}*\rho\simeq_{x,z}\gamma^{\prime}*\rho^{\prime}$.
Finally, the identity arrows are given by the classes of constant paths, while the associativity of composition follows easily with an adaptation of the approach proposed by Brown [4]. Given arrows $[\gamma]\colon w\to x$, $[\gamma^{\prime}]\colon x\to y$ and $[\gamma^{\prime\prime}]\colon y\to z$, let $\Gamma\colon[0,3]\to X$ be the continuous pasting of the representatives of the classes without reparametrizations, i.e.,
$\Gamma(t)\coloneqq\begin{cases}\gamma(t)&\text{ if }t\in[0,1]\\ \gamma^{\prime}(t-1)&\text{ if }t\in[1,2]\\ \gamma^{\prime\prime}(t-2)&\text{ if }t\in[2,3]\end{cases}$
and let $p_{0},p_{1}\colon[0,1]\to[0,3]$ be the homeomorphisms such that $\Gamma\circ p_{0}=\gamma*(\gamma^{\prime}*\gamma^{\prime\prime})$ and $\Gamma\circ p_{1}=(\gamma*\gamma^{\prime})*\gamma^{\prime\prime}$. Since there is a path $H\colon I\to\operatorname{C}\left(I,[0,3]\right)$ for which $H(0)=p_{0}$, $H(1)=p_{1}$ and $H(t)\in\mathbb{P}[0,3]$ for each $t$, it follows that
$\begin{split}\Phi\colon I&\to\operatorname{C}\left(I,X\right)\\ t&\mapsto\Gamma\circ H(t)\end{split}$
is a map witnessing $\gamma*(\gamma^{\prime}*\gamma^{\prime\prime})\simeq_{w,z}(\gamma*\gamma^{\prime})*\gamma^{\prime\prime}$, since the composition is a continuous map under continuous convergence, i.e.,
Proposition 3.1.
The map
$\begin{split}\circ\colon\operatorname{C}\left(X,Y\right)\times\operatorname{C}\left(Y,Z\right)&\to\operatorname{C}\left(X,Z\right)\\ \left\langle g,f\right\rangle&\mapsto f\circ g\end{split}$
is continuous for every limit spaces $X$, $Y$ and $Z$.
Proof.
We present a proof for the sake of completeness: if $\left\langle g_{d},f_{d}\right\rangle_{d}$ converges to $\left\langle g,f\right\rangle$ in $\operatorname{C}\left(X,Y\right)\times\operatorname{C}\left(Y,Z\right)$ and $x_{a}\to x$ in $X$, then $g_{d}(x_{a})\to g(x)$ in $Y$ and $f_{d^{\prime}}(g_{d}(x_{a}))\to f(g(x))$ in $Z$, hence the result follows as $\left\langle f_{d}(g_{d}(x_{a}))\right\rangle_{d,a}$ is a subnet of $\left\langle f_{d^{\prime}}(g_{d}(x_{a}))\right\rangle_{d^{\prime},d,a}$. ∎
With similar reasoning it can be showed that every arrow in $\Pi(X)$ is an isomorphism, so it is in fact a groupoid. Just as happens with the topological fundamental groupoid, the construction above induces a functor $\Pi\colon\textsc{Lim}\to\textsc{Groupoid}$ which sends a continuous map $f\colon X\to Y$ to the functor
$\begin{split}\Pi(f)\colon\Pi(X)&\to\Pi(Y)\\ [\gamma]&\mapsto[f\circ\gamma]\end{split},$
which is well defined because the correspondence $t\mapsto f\circ H(t)$ defines a path in $\operatorname{C}\left(I,Y\right)$ from $f\circ\gamma$ to $f\circ\gamma^{\prime}$ whenever $H\colon I\to\operatorname{C}\left(I,X\right)$ is a path from $\gamma$ to $\gamma^{\prime}$. The details are left to the reader.
By construction, the Lim-groupoid functor above extends the usual Top-groupoid functor $\Pi\colon\textsc{Top}\to\textsc{Groupoid}$, in the sense that the diagram below is commutative [^10] In particular, $\Pi(X)$ is a wide subgroupoid of $\Pi(\tau(X))$ for every limit space $X$..
[…]
Considering the topological modification $\tau\colon\textsc{Lim}\to\textsc{Top}$, one might wonder if $\Pi(X)=\Pi(\tau(X))$ for every limit space $X$. If this were true, our entire approach would seem redundant. However, this is not the case.
Example 3.4.
Over $X\coloneqq\mathbb{R}$, we define a convergence $L$ by putting $\varphi\to_{L}x$ if and only if $\varphi$ is a subnet of a sequence converging to $x$ in ordinary sense. In [10, 11], it is showed that $\left\langle X,L\right\rangle$ is a pseudotopological non-topological space such that $\tau(X)=\mathbb{R}$, i.e., the topological modification of $L$ is the standard topological convergence of the real line. Here we show that $\Pi(X)$ is discrete in the categorical sense, hence $\Pi(X)\neq\Pi(\tau(X))$. First, we need a lemma, whose proof we left to the reader.
Lemma 3.5.
If $\rho\colon I\to\mathbb{R}$ is a non-constant path, then there is $t\in I$ such that $\rho[V]$ is uncountable for every neighborhood $V\subseteq I$ of $t$.
[…]
To obtain the fundamental group of a pointed limit space $(X,x_{0})$, we simply put $\pi_{1}(X,x_{0})\coloneqq\Pi(X)[x_{0},x_{0}]$, as this coincides precisely with the classical definition (cf. Kammeyer [17]). It follows that $\pi_{1}(X,x_{0})$ is a subgroup of $\pi_{1}(\tau(X),x_{0})$ for every pointed limit space $(X,x_{0})$.
Example 3.6.
The previous example yields a non-topological space with trivial fundamental group [^11] For another easy example, recall that for a Tychonoff space $X$, $\operatorname{C}\left(X,\mathbb{R}\right)$ has a topology inducing the continuous convergence if and only if $X$ is locally compact [11]. So, $\left\langle\operatorname{C}\left(X,\mathbb{R}\right),\to_{c}\right\rangle$ is a non-topological limit vector space whenever $X$ is a Tychonoff and non-locally compact space, hence it is null-homotopic.. To get a non-topological space with non-trivial fundamental group, we adapt Example 1.3.2 in [1]. Let $\nu$ be the usual topology of $\mathbb{R}^{2}$ and $X\subseteq\mathbb{R}^{2}$ be a “lollipop”, i.e., the union of a line $L$ and a circle $C$ of the plane, such that $L\cap C=\{p\}$, as in Figure 1 below.
Figure 1. The space $X$.
Let $S\coloneqq L\setminus\{p\}$ and $D\subseteq S$ be a countable $\nu$-dense subset of $S$. For a net $\varphi\in\textsc{NETS}\left(X\right)$ and a point $x\in X$, we define a limit convergence $\lambda$ such as:
  1. (1)
    in case $x\neq p$, then $\varphi\to_{\lambda}x$ if and only if $\varphi\to_{\nu}x$;
  2. (2)
    in case $x=p$, then $\varphi\to_{\lambda}x$ if and only if
    1. (a)
      $C\in\varphi^{\uparrow}$ and $\varphi\to_{\nu}p$, i.e., there is $A\in\operatorname{dom}\left(\varphi\right)$ such that $\varphi(a)\in C$ for every $a\succeq A$ and $\varphi\to_{\nu}p$ in the usual sense, or
    2. (b)
      $\varphi^{\uparrow}\mathrel{\#}D$ and $\varphi\to_{\nu}p$, i.e., for every $A\in\operatorname{dom}\left(\varphi\right)$ there is $a\succeq A$ such that $\varphi(a)\in D$ and $\varphi\to_{\nu}p$ in the usual sense, or
    3. (c)
      $\varphi$ is a subnet of a mixing of finitely many nets satisfying conditions (a) and (b).
The definition of $\lambda$ ensures $\left\langle X,\lambda\right\rangle\in\textsc{Lim}$. To see that $\left\langle X,\lambda\right\rangle\not\in\textsc{Top}$, notice that $\operatorname{adh}_{\lambda}\left(A\setminus C\right)=A\setminus\{p\}$ and $\operatorname{adh}_{\lambda}\left(A\setminus\{p\}\right)=A$, thus showing that the adherence operator is not idempotent. It is not hard to see that $\pi_{1}(X,p)=\pi_{1}(C,p)$: since the topological modification of $\left\langle X,L\right\rangle$ is the $\nu$-subspace topology inherited from $\mathbb{R}^{2}$, it follows that a continuous path $\gamma\colon I\to X$ such that $\gamma(0)=p$ cannot intercept $S$, i.e., there is no $t\in I$ such that $\gamma(t)\in S$, otherwise we would be able to obtain a sequence $\left\langle x_{n}\right\rangle_{n}$ in $I$ such that $x_{n}\to 0$ but $\gamma(x_{n})\in A\setminus C$ for every $n$.
4. The Seifert-Van Kampen Theorem for fundamental groupoids of limit spaces
Let us recall one of the many versions of the theorem in the title, as the one presented in [17].
Theorem 4.1 (Seifert-Van Kampen, groupoid version).
Let $X$ be a topological space, and let $\mathcal{O}$ be an open cover of $X$, which is closed under finite intersections. Consider $\mathcal{O}$ as a small category with morphisms given by inclusions. Then, restricting $\Pi$ to $\mathcal{O}$ defines a diagram $\Pi|_{\mathcal{O}}\colon\mathcal{O}\to\textsc{Groupoid}$ such that $\Pi(X)=\operatorname{colim}\left(\Pi|_{\mathcal{O}}\right)$.
The purpose of this section is to replace the word “topological space” with “limit space” in a reasonable way. To do that, all we have to do is to replace “open covers” with “convergence systems”, as defined in Section 2, and carry over the usual proof with the obvious adaptations.
Theorem 4.2 (Seifert-Van Kampen, groupoid version in Lim).
Let $X$ be a limit space, and let $\mathcal{O}$ be a convergence system for $X$, which is closed under finite intersections. Consider $\mathcal{O}$ as a small category with morphisms given by inclusions. Then, restricting $\Pi$ to $\mathcal{O}$ defines a diagram $\Pi|_{\mathcal{O}}\colon\mathcal{O}\to\textsc{Groupoid}$ such that
$\Pi(X)=\operatorname{colim}\left(\Pi|_{\mathcal{O}}\right).$
Sketch of the proof.
We follow [17]. The idea is to show that $\left\langle\Pi(U)\to\Pi(X)\right\rangle_{U\in\mathcal{O}}$ is a universal cocone in the category of cocones of $\Pi|_{\mathcal{O}}$. As it is clearly a cocone, we consider any other cocone $\left\langle\mathcal{F}_{U}\colon\Pi(U)\to\mathcal{G}\right\rangle_{U\in\mathcal{O}}$, where $\mathcal{G}$ is a groupoid, and show that there exists a unique functor $\mathcal{F}\colon\Pi(X)\to\mathcal{G}$ such that
${{\Pi(U)}}$${{\mathcal{G}}}$${{\Pi(X)}}$$\scriptstyle{\mathcal{F}_{U}}$$\scriptstyle{\mathcal{F}}$
commutes for every $U\in\mathcal{O}$. On objects, we put $\mathcal{F}(x)\coloneqq\mathcal{F}_{U}(x)$ for any $U$ such that $x\in U$. Such a $U$ exists because the convergence on $X$ is centered, and it is well-defined since $\left\langle\mathcal{F}_{U}\right\rangle_{U\in\mathcal{O}}$ is a cocone and $\mathcal{O}$ is closed under finite intersections.
To define $\mathcal{F}([\gamma])$ for an arrow $[\gamma]\colon x\to y$, we use a Lebesgue-$\delta$ number for the convergence system $\mathcal{O}$ (cf. Proposition 2.1). For sufficiently large $n$, the interval $I$ can be subdivided into $n$ subintervals of length less than $\delta$. Hence we can write $\gamma=\gamma_{1}*\dotso*\gamma_{n}$, where each $\gamma_{i}$ is a path in $X$ whose image is contained in some $U_{i}\in\mathcal{O}$. Finally, set $\mathcal{F}([\gamma])=\mathcal{F}_{U_{n}}([\gamma_{n}])\circ\dotso\circ\mathcal{F}_{U_{1}}([\gamma_{1}])$. The arguments that ensure $\mathcal{F}$ is well-defined and functorial are similar to those in the topological case. ∎
Example 4.3.
Back to Example 3.4, notice that for a point $x_{0}\in X$, the family $\mathcal{N}\coloneqq\{N\cup\{x_{0}\}:N\subseteq X$ and $N$ is countable}}\}} is a convergence system for $X$, closed under finite intersections, so $\Pi(X)=\operatorname{colim}_{N^{\prime}\in\mathcal{N}}(\Pi(N^{\prime}))$. Now, as each $N^{\prime}\in\mathcal{N}$ is countable, it follows easily that every continuous function $\gamma\colon I\to N^{\prime}$ is constant, otherwise $\tau(\gamma)\colon I\to\tau(N^{\prime})\subseteq\mathbb{R}$ would be a non-constant continuous function, which is absurd as $\tau(N^{\prime})$ has $N^{\prime}$ as its underlying set, which is too small to contain a non-degenerated open interval of the real line.
5. Further remarks and comments
The results presented in this work illustrate, to some extent, the broad scope of investigation within this program of “extending” Algebraic Topology towards what might be called “Algebraic Convergence”. A natural next step would be to extend or adapt results concerning universal coverings to convergence spaces. This was recently proposed by Treviño-Marroquín in [37], though using the language of filters [^12] Interestingly, the definition of connectedness used in [37] is equivalent to the one presented in [11], which essentially requires that the topologization of the convergence space be a connected topological space.. Even more recently, Milićević and Scoville  [23] discussed how to further develope singular homology and higher homotopy groups in the category of pseudotopological spaces.
In another direction, one could argue that the generality of limit spaces calls for a less restrictive notion of fundamental group(oid). With this in mind, and considering the alternatives presented by Kennison in [20], a natural question to do is the following: What is the correct generalization of a sheaf for limit spaces? Could this be used to describe fundamental groups of limit spaces?
Acknowledgments
We would like to thank Germán Ferrer, Peter Wong and Weslem Silva for their insightful and helpful discussions on the subject.